Solve y zp + x²zq = y2x
1
Expert's answer
2020-10-17T14:18:28-0400

"\displaystyle\textsf{The given partial differential equation is}\\

 yz\, p + x^2 z\, q = y^2 x \\


\mathrm{d}z =\frac{\partial z}{\partial x} \mathrm{d}x + \frac{\partial z}{\partial y} \mathrm{d}y \\

\mathrm{d}z = xy^2, \mathrm{d}x = yz, \mathrm{d}y = x^2 z \\


\textsf{The Lagrange’s auxiliary equations}\\\textsf{for the given PDE is}\\

\frac{\mathrm{d}x}{yz} = \frac{\mathrm{d}y}{x^2 z} = \frac{\mathrm{d}z}{xy^2}\\

\textsf{Choosing}\, (x^2, -y, 0)\, \textsf{as multipliers, we have}\\

x^2\mathrm{d}x - y\mathrm{d}y = 0 \\

\textsf{Integrating both sides, we have} \\

\int x^2\mathrm{d}x - \int y\mathrm{d}y = C_1 \\

\frac{x^3}{3} - \frac{y^2}{2} =c \\

\textsf{Comparing the first and the last differential equations}\\

\frac{\mathrm{d}x}{yz} = \frac{\mathrm{d}z}{y^2 x}\\

yx\mathrm{d}x - z\mathrm{d}z = 0\\


\textsf{But}\\

x^3/3 - y^2/2 = c \\

y = \sqrt{\frac{2x^3}{3} + C} \\

\therefore \int x\sqrt{\frac{2x^3}{3} + C}\, \mathrm{d}x - \frac{z^2}{2} = C_2\\


\textsf{The Integral cannot be expressed in simpler}\\ \textsf{terms and as such it can be expressed}\\\textsf{in terms of advanced functions like the}\\\textsf{hypergeometric function. We can hence}\\\textsf{write the Integral in terms of know}\\\textsf{functions if}\,C = 0.\\

\therefore \int x\sqrt{\frac{2x^3}{3}}\, \mathrm{d}x - \frac{z^2}{2} = C_2\\

\frac{2}{7}\sqrt{\frac{2}{3}} x^{\frac{7}{2}} - \frac{z^2}{2} = C_2\\


\textsf{So, a solution of the given PDE is}\, \\ \phi\left(\frac{x^3}{3} - \frac{y^2}{2}, \frac{2}{7}\sqrt{\frac{2}{3}} x^{\frac{7}{2}} - \frac{z^2}{2}\right) = 0"